OpenStudy (anonymous):

Two integers are relatively prime if they have no common factors other that 1 or -1. What is the probability that a positive integer less than or equal to 30 is relatively prime to 30? Express your answer as a common fraction. Show work

OpenStudy (nikvist):

1,7,11,13,17,19,23,29 (total 8 numbers) p=8/30=4/15